OpenStudy (anonymous):

Grades can be converted into numbers ( as in calculating a GPA where an A=4, B=3,C=2, D=1, and F=0 x Grade 4 3 2 1 0 p(X) 0.20 0.25 0.30 0.10 0.15 a. verify that this is a legitimate probability distribution 0.20+0.25+0.30+0.10+0.15=1 b.Find the means grade for this class. 4X0.20=0.8 3X0.25=0.75 2X0.30=0.6 1X0.10=0.1 0X0.15=0.0 0.8+0.75+0.6+0.1+0.0=2.25 Find the standard deviation c. 2.25/5=0.45

OpenStudy (anonymous):

What do you have to do for part (c)?

OpenStudy (anonymous):

Find the standard deviation

OpenStudy (anonymous):

Okay, so the std dev is the square root of the variance, or \[s=\sqrt{s^2}\] A particularly useful formula for finding the variance is \[V(X)=E(X^2)-\bigg[E(X)\bigg]^2\] In other terms, \[s^2=\sum_{\text{all }x}x^2p(x)-\left(\sum_{\text{all }x}x~p(x)\right)^2\] Then for the std dev, you would just take the square root.

OpenStudy (anonymous):

how did get 8

OpenStudy (anonymous):

Given this data, you have \[s^2=\bigg(4^2\cdot0.20+3^2\cdot0.25+\cdots+0^2\cdot0.15\bigg)-\bigg(2.25\bigg)^2\]

OpenStudy (anonymous):

go the same ans but did it different

OpenStudy (anonymous):

is that how I get c

OpenStudy (anonymous):

c=1.68 or 1.69